The circle has diameter RS where R is (-3,2) and S is (3,2). Find the equation of the circle.
1
Expert's answer
2011-05-17T11:14:20-0400
Let's find the center of the circle, it lays on the middle of the interval RS and has the coordinates:
(3+(-3)/2, (2+2)/2)
O (0, 2).

The radius of this circle would be
R2 = (3-0)2 + (2-2)2 = 9
R = 3.

Thus the equation of the circle is
x2 + (y-2)2 = 9